OpenStudy (anonymous):

Find the perimeter of a rectangle with l = 3.2 m and w = 0.5 m. A. 1.6 m B. 3.2 m C. 3.7 m D. 7.4 m

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

2L + 2W = perimeter of the rectangle?..

OpenStudy (anonymous):

or letter D instead :D

OpenStudy (mathmale):

The equation for the perimeter of a rectangle of length L and width W is, as Shikatonori has already typed in, is P = 2L + 2W. You are given the values of L and W. Substitute them into this equation for the perimeter.

OpenStudy (anonymous):

What is the perimeter of the polygon? A. 93.3 cm B. 103.3 cm C. 103.5 cm D. 104.3 cm

OpenStudy (anonymous):

@Shikatanori

OpenStudy (freckles):

did you try pluggin into @mathmale 's formula that he provided ?

OpenStudy (anonymous):

no

OpenStudy (freckles):

P=2L+2W replace L with 3.2 and W with 0.5

OpenStudy (anonymous):

What is the perimeter of the polygon? A. 93.3 cm B. 103.3 cm C. 103.5 cm D. 104.3 cm

OpenStudy (freckles):

you just need to multiply a couple of pairs of numbers and then add their products you can do it

OpenStudy (freckles):

or if don't like to multiply instead of looking at 2L+2W do L+L+W+W

OpenStudy (freckles):

if your asking about a different figure i don't see one that has been provided

OpenStudy (anonymous):

is 103.5

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

OpenStudy (anonymous):

Unit 2 | Lesson 3 Lesson Checkpoint: Show What You Know Question Navigator 5. Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

OpenStudy (anonymous):

Well P = 2L + 2W = 2 . 24 +2 . 18

OpenStudy (anonymous):

Just Add all the sides of the polygon.. Done simply like that!... :) its the peremeter!! @coolman54